OpenStudy (anonymous):

Find derivative of t-4sqrt(t+1) I got 1-2(t+1)^(-1/2)

OpenStudy (abb0t):

well, do you know that \(\sqrt{t+1} = (t+1)^{\frac{1}{2}}\) that might help

OpenStudy (anonymous):

yeah that I knew, I know that the derivative of t=1, and 4sqrt(t+1) is just 4(t+1)^(1/2)

OpenStudy (anonymous):

now the derivative of 4(t+1)^(1/2) is just 2(1+x)^(-1/2)